# T3-S2S: Training-free Triplet Tuning for Sketch to Scene Generation

### v1

@@ 6 Conclusion
Appendix A Discussion
While our approach is innovative and enhances multi-instance scene generation, it also has some room to improve, primarily stemming from the inherent capabilities of the base model. One significant challenge is the generation of detailed instances, such as textures and finer details. This issue largely arises from the limited understanding of complex descriptions by the CLIP models. Moreover, the characteristic prominence module tends to focus on instance tokens while neglecting some descriptive adjectives. Additionally, our method struggles with accurately capturing very large scenes (exceeding $4096\times 4096$ pixels) such as expansive game maps, which often include complex relationships like overlaps and interactions between instances. These complex and dynamic scenarios require further enhancements and refinements in our approach to effectively represent and capture such intricate relationships. Building on our current achievements, we plan to further explore these areas in future work to improve detailed multi-instance sketch-to-scene generation.
Appendix B Top $K$ Anaylsis.
(a) The first and second extreme points.
(b) The third and fourth extreme points.
Figure 9: Histogram of the distribution of indices where the extremum points are located.
In the module of characteristic prominence, two hyperparameters, $K$ and $\beta$, require meticulous tuning. $K$ determines the indices of extreme values within the value matrices. For our analysis, we save these indices and construct a histogram, as depicted in Figure 9. We observe that the probability of later instance tokens achieving the maximum and second maximum values is comparatively lower than that of earlier tokens. Thus, increasing the value of later instance tokens will be beneficial for their representations. However, at the third and fourth extremes, the probabilities tend to converge, indicating that not every token is essential for defining key characteristics. Increasing values for later instances at this point would introduce additional noise. Therefore, setting $K=2$ is advisable based on the observed trends. For $\beta$, which enhances the characteristics of instances within the feature matrix, an initial increase is beneficial. Nonetheless, there is a critical threshold beyond which further increases in $\beta$ begin to disrupt the distribution within the value matrices.
Appendix C Sketch Visualizations of Quantitative Experiments
@@
Using a game scene as an example, we begin each prompt with ‘Isometric view of a game scene’ to generate controlled synthetic images for game settings. This helps maintain a consistent angle and style, ignoring any incoherent instance sketches that might appear in real-world scenes, thereby focusing on object placement and verifying text-image consistency. We generate all 20 complex scenes using hyperparameters identical to those used in the main results (Figure 6), shown in Figures 12 and 13. The colored sketches are used solely to distinguish between different instances, and the colors used are arbitrary without class or semantic information. To validate this, we also use grayscale sketches as input, and the resulting images are nearly identical under the same random seed (two columns pointed by the red arrows in Figure 12). Meanwhile, our approach is not limited to game scenes. We also test prompts without the fixed game scene phrase, resulting in more diverse angles and styles while maintaining the same quality in object placement and text-image consistency (One row pointed by the green arrows in Figures 12).
Appendix D Visualization of Diverse Scenes
@@
In the above experiments, we primarily validate the controllability of our method for multi-instance generation in game scenes. However, this does not imply that our approach is limited to game scenarios. To further verify its capabilities, we design three sets of diverse scenes: (1) four common simple scenes; (2) two indoor scenes; and (3) three scenes featuring instances of the same type but with different color attributes. Without changing any hyperparameters, generations are presented in Figure 14. In common scenes, our method effectively mitigates instance overlap under ControlNet control, while in indoor scenes, it handles varied layouts well. For the challenging task of differentiating attributes within identical instances, our approach assigns distinct properties accurately. However, for uncommon attributes like generating a red cat, our method struggles due to limitations inherent in the original SDXL model.
@@ Appendix E Metric of User Study
We conduct a user study on 20 scenes, each with 6 variants, generating 100 images per scene. A Gradio-based evaluation interface is designed, which randomly selects one image from 120 sets to create a sub-evaluation system, with images presented anonymously. 23 participants independently rate the images based on the following scale:
@@
Appendix F Transfer Prompt Balance to Attend-and-Excite
Figure 10: Visualizations for transferring PB to Attend-and-Excite [42]. In most cases, both instances are successfully generated. The frog-leg cat and the bird-wing pig further demonstrate the effectiveness since they lack the layouts to separate the instances spatially.
To further validate the PB module, we integrated it into the Attend-and-Excite method [42], based on attention tuning using the SD V1.4 model. The results are shown in Figure 10. Despite the limitations of SD V1.4, the PB module effectively balances embedding strength between the two instances in scenarios without layout guidance, enhancing their representation. In most cases, both instances are successfully generated. However, in some cases, the attributes of the two objects become entangled, leading to artifacts such as a cat with frog legs or a pig with bird wings, due to the lack of spatial separation, which further demonstrates the effectiveness of the PB module.
@@ Appendix G Transfer T3-S2S to T2I-Adapter
Figure 11: Visualizations for transferring T3-S2S to T2I-Adapter [30]. T3-S2S effectively improves the T2I-Adapter’s alignment with prompts and layouts in complex scenes, demonstrating its control capabilities across different models.
@@
To validate the general applicability of our approach beyond the ControlNet model, we apply T3-S2S to another controllable T2I-Adapter [30] model. Although the T2I-Adapter performs best with detailed sketches, we use grayscale sketches for quick validation, which contain less semantic information. We keep the PB and CP modules unchanged, while the DT module is integrated into the SDXL main channel, similar to CP, as it can not be placed in a separate branch like in ControlNet. We use the same prompts and sketches from the main results (Figure 6) and Appendix C, with all other hyperparameters unchanged. The results are shown in Figure 11. T3-S2S effectively improves the T2I-Adapter’s alignment with prompts and layouts in complex scenes, demonstrating its control capabilities across different models. However, the generation quality still lags behind the ControlNet-based approach, indicating the need for parameter tuning specific to the T2I-Adapter’s distribution and improved sketch inputs to align with the T2I-Adapter. Despite these limitations, the results show that T3-S2S has promising generalizability and can effectively control both ControlNet and T2I-Adapter models.
Appendix H 3D Game Scene
Figure 15 demonstrates examples of 3D scenes generated using our method. Building on the approach in  [33], our method can be used to reconstruct a 3D mesh and further serve as the foundation for generating high-fidelity 3D scenes within the game environment. Similar to  [33], we also adopt the Depth-Anything-V2 [43] method to infer scene depth and reconstruct the complete mesh using the Poisson reconstruction technique.
@@
Figure 12: Example results from a subset of the 20 complex scene composition tested using hyperparameters identical to those used in the main results (Figure 6). (1) Two columns pointed by the red arrows represent the generations using colored and grayscale sketches under the same random seed. (2) One row pointed by the green arrows indicates the generations without the fixed game scene phrase.
@@
Figure 13: Example results from a subset of the 20 complex scene composition tested using hyperparameters identical to those used in the main results (Figure 6). Two columns pointed by the red arrows represent the generations using colored and grayscale sketches under the same random seed.
@@
Figure 15: Example of 3D scene generation results. The left side displays the input sketches and text, along with the generated isometric images. The images on the right are rendered from the reconstructed 3D scene using the isometric images.

### v2

@@ 6 Conclusion
Appendix A Discussion and Limitations
While our approach is innovative and enhances multi-instance scene generation, it also has some room to improve, primarily stemming from the inherent capabilities of the base model.
Texture Generation. Generating fine-grained textures and detailed instances remains challenging. This issue is partly due to the limited ability of CLIP-based representations to precisely encode complex descriptive attributes. Furthermore, the characteristic prominence module prioritizes instance-level tokens, which can lead to reduced attention on adjective-level descriptors that are critical for texture and appearance modeling. For example, as illustrated in Fig. 17, in cases where multiple instances of the same object with different colors are specified, our method is generally able to generate the desired targets. However, attribute confusion may still occur, leading to mismatches between instances and their corresponding color attributes.
Large Scene Challenge. Our method encounters difficulties in accurately generating expansive scenes (e.g., game maps exceeding $4096\times 4096$ pixels). Processing such high resolutions directly incurs prohibitive computational costs, typically necessitating patch-based processing or upsampling strategies. However, a more fundamental constraint lies in the model architecture itself. Taking SDXL as an example, the network downsamples a $1024\times 1024$ input to a minimum bottleneck feature size of $32\times 32$. This imposes an intrinsic $32\times$ scaling limit, meaning the minimum controllable granularity for object generation is restricted to this ratio. For example, in the first case of Figure 16, the control over minute objects degrades at the bottleneck layer. Consequently, even with increased output resolution, accurately resolving and controlling tiny instances or intricate interactions below this structural floor remains a significant challenge.
Concept Art Understanding. Additionally, the 3D scene shows that terrain and objects can be generated as concept art, but improvements are needed in understanding concept art, particularly in segmentation and depth estimation, which may not be accurately represented in the synthesized data. One potential remedy is to concurrently generate multiple modalities at the same time, such as RGB, semantic, depth, material, and object footprint, and fuse these results.
Building on our current achievements, we plan to further explore these areas in future work to improve detailed multi-instance sketch-to-scene 3D generation.
Appendix B Details of Dense Tuning
Dense Diffusion (Kim et al., 2023) can effectively amplify instance-specific attention responses in standard diffusion models. However, when adapting dense modulation strategies to ControlNet, it is crucial to account for the hierarchical roles of different layers. We follow the layer-and-step-wise empirical observations (Voynov et al., 2023; Wang et al., 2024a; Sun et al., 2024) that peripheral (early) layers predominantly encode low-level geometric cues and boundary information, while more central layers are responsible for shaping object contours and structural layouts. This behavior is consistent with the hierarchical design of diffusion UNets and ControlNet, where early steps preserve fine spatial details and later steps increasingly capture semantic structure. As demonstrated in Table 2, DT alone yields marginal quantitative improvement, indicating that DT primarily acts as a modulation mechanism rather than the main source of performance gain, which is instead driven by the proposed PB and CT modules. Thus, to evaluate its application, we introduce adapted variants of DT across different layers within our proposed PB+CT framework.
Qualitative analysis. Figure 12 presents a qualitative comparison of Dense Tuning (DT) applied at different depths of ControlNet. Applying DT at mid_block_0 already produces visually coherent and well-controlled results, while extending DT to down_block_2 + mid_block_0 further improves structural consistency. In contrast, including earlier layers such as down_block_1 introduces overly strong contour and edge responses, resulting in visually intrusive ControlNet outlines that degrade image naturalness. This observation indicates that early ControlNet layers predominantly encode low-level geometric cues, where dense modulation tends to over-amplify sketch boundaries rather than semantic structure.
Table 3: Cost–benefit analysis of Dense Tuning (DT) applied to different layer combinations in ControlNet and U-Net within the PB+CT framework.
[…]
M0 N/A 18175 10.20 0.3453 0.2532 0.2548
D2 + M0 N/A 18215 10.61 0.3491 0.2566 0.2581
D1 + D2 + M0 N/A 18498 10.89 0.3429 0.2501 0.2522
D2 + M0 M0 18661 10.96 0.3494 0.2571 0.2569
D2 + M0 D2 + M0 + U0 18949 11.14 0.3411 0.2492 0.2552
D2 + M0 D1 + D2 + M0 + U0 + U1 19297 11.35 0.3401 0.2496 0.2547
Figure 12: Qualitative comparison of Dense Tuning (DT) applied to different ControlNet layers. DT at mid_block_0 yields coherent and well-controlled results, while extending DT to down_block_2 + mid_block_0 further improves structural consistency. In contrast, applying DT to earlier layers (e.g., down_block_1) over-amplifies sketch contours and edges, leading to visually intrusive ControlNet outlines and degraded image naturalness. Applying DT to mid-level U-Net layers leads to marginal visual differences, whereas tuning peripheral U-Net layers introduces noticeable degrading changes. Prompts and Sketches are from Figures 915, and 16.
Efficiency analysis. Table 3 reports GPU memory consumption and inference latency under different DT layer configurations. As DT is activated in additional layers, both memory usage and inference time increase consistently. Given the limited performance gains observed in Table 2 and the growing computational overhead shown in Table Y, restricting DT to down_block_2 and mid_block_0 yields the most favorable trade-off between controllability, efficiency, and overall visual quality.
Based on these qualitative and quantitative observations, we adopt down_block_2 + mid_block_0 as the default DT configuration in ControlNet.
Appendix C Visualization of Game Scenes
@@
Using a game scene as an example, we begin each prompt with ‘Isometric view of a game scene’ to generate controlled synthetic images for game settings. This helps maintain a consistent angle and style, ignoring any incoherent instance sketches that might appear in real-world scenes, thereby focusing on object placement and verifying text-image consistency. We generate all 50 complex scenes using hyperparameters identical to those used in the main results (Fig. 7 in the main paper), shown in Fig. 12 in the main paper and Fig. 16. The colored sketches are used solely to distinguish between different instances, and the colors used are arbitrary without class or semantic information. To validate this, we also use grayscale sketches as input, and the resulting images are nearly identical under the same random seed (two columns pointed by the red arrows in Fig. 16). Meanwhile, our approach is not limited to game scenes. We also test prompts without the fixed game scene phrase, resulting in more diverse angles and styles while maintaining the same quality in object placement and text-image consistency (One row pointed by the green arrows in Fig. 12 in the main paper.
Appendix D Visualization of Common Scenes
@@
Qualitative Common Scenes. In the above experiments, we primarily validate the controllability of our method for multi-instance generation in game scenes. However, this does not imply that our approach is limited to game scenarios. To further verify its capabilities, we design three sets of diverse scenes: (1) four common simple scenes; (2) two indoor scenes; and (3) three scenes featuring instances of the same type but with different color attributes. Without changing any hyperparameters, generations are presented in Fig. 17. In common scenes, our method effectively mitigates instance overlap under ControlNet control, while in indoor scenes, it handles varied layouts well. For the challenging task of differentiating attributes within identical instances, our approach assigns distinct properties accurately. However, for uncommon attributes like generating a red cat, our method struggles due to limitations inherent in the original SDXL model.
@@ Appendix E Metric of User Study
We conduct a user study on 50 scenes, each with 6 variants, generating 100 images per scene. A Gradio-based evaluation interface is designed, which randomly selects one image from 120 sets to create a sub-evaluation system, with images presented anonymously. 23 participants independently rate the images based on the following scale:
@@
Appendix F Transfer PB to Attend-and-Excite
Figure 13: Visualizations for transferring PB to Attend-and-Excite (Chefer et al., 2023). In most cases, both instances are successfully generated. The frog-leg cat and the bird-wing pig further demonstrate the effectiveness since they lack the layouts to separate the instances spatially.
To further validate the PB module, we integrated it into the Attend-and-Excite method (Chefer et al., 2023), based on attention tuning using the SD V1.4 model. The results are shown in Figure 13. Despite the limitations of SD V1.4, the PB module effectively balances embedding strength between the two instances in scenarios without layout guidance, enhancing their representation. In most cases, both instances are successfully generated. However, in some cases, the attributes of the two objects become entangled, leading to artifacts such as a cat with frog legs or a pig with bird wings, due to the lack of spatial separation, which further demonstrates the effectiveness of the PB module.
@@ Appendix G Transfer T3-S2S to T2I-Adapter
Figure 14: Visualizations for transferring T3-S2S to T2I-Adapter (Mou et al., 2023). T3-S2S effectively improves the T2I-Adapter’s alignment with prompts and layouts in complex scenes, demonstrating its control capabilities across different models.
@@
To validate the general applicability of our approach beyond the ControlNet model, we apply T3-S2S to another controllable T2I-Adapter (Mou et al., 2023) model. Although the T2I-Adapter performs best with detailed sketches, we use grayscale sketches for quick validation, which contain less semantic information. We keep the PB and CP modules unchanged, while the DT module is integrated into the SDXL main channel, similar to CP, as it can not be placed in a separate branch like in ControlNet. We use the same prompts and sketches from the main results (Fig. 7 in the main paper) and Appendix C, with all other hyperparameters unchanged. The results are shown in Figure 14. T3-S2S effectively improves the T2I-Adapter’s alignment with prompts and layouts in complex scenes, demonstrating its control capabilities across different models. However, the generation quality still lags behind the ControlNet-based approach, indicating the need for parameter tuning specific to the T2I-Adapter’s distribution and improved sketch inputs to align with the T2I-Adapter. Despite these limitations, the results show that T3-S2S has promising generalizability and can effectively control both ControlNet and T2I-Adapter models.
Appendix H Perception Modeling for Structural Assembly
To bridge the transition between twin structural concept representations, i.e., isometric and terrain images, we follow sketch2scene (Xu et al., 2024) as our backend to build 3D game scenes from concept art generations. Structurally Specified Assembly module constructs full 3D natural scenes by segmenting foreground assets from isometric images, generating terrain meshes via tomography from terrain images, and applying sketch-based textures. This supports accurate terrain modeling, rich asset composition, and seamless integration with standard engines.
Foreground Instances. We extract 2D isometric images of individual objects (e.g., buildings, houses, trees) from the instance segmentation map via Segment Anything (Kirillov et al., 2023), which are later used for 3D retrieval or conditioning in the 3D generation. Assuming an isometric projection with a $45^{\circ}$ yaw angle, we estimate object poses by applying homography warping (Hartley and Zisserman, 2003) and analyzing their bounding boxes. The warped instance segmentation provides the object footprint, while the depth information enables projection into 3D space.
HeightMap Generation. The isometric view offers sufficient coverage to recover both geometry and appearance from a single terrain image. We reconstruct a watertight terrain mesh from the image using Depth Anything (Yang et al., 2024) for depth estimation, followed by Poisson surface reconstruction. From the coarse depth, we obtain an isometric view depth map and compute the height as $h=d_{\max}-d$. The height is rotated into bird’s-eye view (BEV) via homography warping (Hartley and Zisserman, 2003) to build the terrain mesh, and then the color-aligned mesh is segmented via Segment Anything (Kirillov et al., 2023) and Osprey (Yuan et al., 2023) into categories (grass, rock, mud, road, etc.). The categories map is used to retrieve standard texture tiles to ensure high-res appearance in close-up views. For water regions, we lower the terrain and insert a water asset to maintain realistic elevation.
Procedural 3D Scene Synthesize. Using semantic and geometric data, we procedurally generate and render 3D scenes via game engines (e.g., Unity, Unreal). We adopt Unity for its terrain, vegetation, and animation support. Heightmaps and splatmaps are mapped to Unity terrain. Vegetation is procedurally placed based on texture semantics (e.g., grass → flowers, rocks). Foreground objects are placed by retrieval (from Objaverse via CLIP) or generated using 2D-to-3D models (Hong et al., 2023; Wei et al., 2024; hyperhuman, 2024), ensuring style consistency with the reference.
3D scene generations following the sketch2scene (Xu et al., 2024) are shown in Figs. 19 and  20.
@@ Appendix H Perception Modeling for Structural Assembly
Figure 15: Example results from a subset of the 50 complex scene compositions tested using hyperparameters identical to those used in the main results (Fig. 9). (1) Two columns pointed by the blue arrows represent the structural concept representations using the prefix terrain prompts and isometric prompts under the same random seed. (2) Two columns pointed by the red arrows represent the generations using colored and grayscale sketches under the same random seed. (3) One row pointed by the green arrows indicates the generations without the fixed game scene phrase.
@@
Figure 16: Example results from a subset of the 50 complex scene compositions tested using hyperparameters identical to those used in the main results (Fig. 7). (1) Two columns pointed by the blue arrows represent the structural concept representations using the prefix terrain prompts and isometric prompts under the same random seed. (1) Two columns pointed by the red arrows represent the generations using colored and grayscale sketches under the same random seed.
@@
Figure 18: Another visual comparison of two hyper-parameters $K$ and $\beta$ suggests that setting $K=2$ and $\beta=1$ is a favorable choice. With $\beta=1$, increasing $K$ initially improves results but eventually adds noise. Fixing $K=2$, higher $\beta$ values yield stable and favorable outcomes.
Figure 19: 3D scene generation results with two examples from Fig. 9. The 3D scene shows that terrain and objects can be generated as concept art, but improvements are needed in understanding concept art, particularly in segmentation and depth estimation, which may not be accurately represented in the synthesized data. Additional 3D scenes are available in the Appendix and the supplementary video materials.
Figure 20: 3D scene generation results with two more examples from Fig. 16. The 3D scene shows that terrain and objects can be generated as concept art, but improvements are needed in understanding concept art, particularly in segmentation and depth estimation, which may not be accurately represented in the synthesized data.